# envname-P envname#1

### v1

@@ Abstract
Flow models parameterized as time-dependent velocity fields can generate data from noise by integrating an ODE. These models are often trained using flow matching, i.e. by sampling random pairs of noise and target points $(\mathbf{x}_{0},\mathbf{x}_{1})$ and ensuring that the velocity field is aligned, on average, with $\mathbf{x}_{1}-\mathbf{x}_{0}$ when evaluated along a time-indexed segment linking $\mathbf{x}_{0}$ to $\mathbf{x}_{1}$. While these noise/data pairs are sampled independently by default, they can also be selected more carefully by matching batches of $n$ noise to $n$ target points using an optimal transport (OT) solver. Although promising in theory, the OT flow matching (OT-FM) approach (Pooladian et al., 2023; Tong et al., 2024) is not widely used in practice. Zhang et al. (2025) pointed out recently that OT-FM truly starts paying off when the batch size $n$ grows significantly, which only a multi-GPU implementation of the Sinkhorn algorithm can handle. Unfortunately, the pre-compute costs of running Sinkhorn can quickly balloon, requiring $O(n^{2}/\varepsilon^{2})$ operations for every $n$ pairs used to fit the velocity field, where $\varepsilon$ is a regularization parameter that should be typically small to yield better results. To fulfill the theoretical promises of OT-FM, we propose to move away from batch-OT and rely instead on a semidiscrete formulation that can leverage the fact that the target dataset distribution is usually of finite size $N$. The SD-OT problem is solved by estimating a dual potential vector of size $N$ using SGD; using that vector, freshly sampled noise vectors at train time can then be matched with data points at the cost of a maximum inner product search (MIPS) over the dataset. Semidiscrete FM (SD-FM) removes the quadratic dependency on $n/\varepsilon$ that bottlenecks OT-FM. SD-FM beats both FM and OT-FM on all training metrics and inference budget constraints, across multiple datasets, on unconditional/conditional generation, or when using mean-flow models.
@@ 1 Introduction
Figure 1: I-FM (left) assigns noise to data purely at random. OT-FM (middle-left) samples batches of $n$ noise and $n$ data points and re-aligns them with an optimal matching permutation $\sigma^{\star}$. These matches are, however, inherently unstable, as $n$ points do not reflect the whole noise distribution nor the dataset. Increasing drastically $n$ can mitigate this issue (Zhang et al., 2025), but at a significant cost. Our method, SD-FM (right), solves these issues in two steps: in a precompute phase, the semidiscrete OT problem (parameterized as a vector of size NN, the dataset size) is solved using SGD. At FM train time , each newly sampled noise is assigned to a data point using a maximum inner product search, Laguerre cells (Mérigot, 2011) being illustrated in the plot. Our figure uses no entropic regularization ($\varepsilon=0$) and a neg-dot-product cost for simplicity, see equation 2.4 for more generality.
@@ 2 Background and related work
Access to $\bm{v}_{t}$ then allows for samples $\bm{X}_{0}\sim\mu_{0}$ to be transformed into samples $\bm{X}_{1}\sim\mu_{1}$ through evolution of equation 2.1. Training a neural approximation $\bm{v}_{\theta}$ for the flow field $\bm{v}$ can be done by sampling $(\bm{X}_{0},\bm{X}_{1})\sim\pi$ and $t\in[0,1]$, setting $\bm{X}_{t}:=\varphi_{t}(\bm{X}_{0},\bm{X}_{1})$ and minimizing
@@
The linear interpolant $\varphi_{t}(\mathbf{x},\mathbf{y}):=(1-t)\mathbf{x}+t\mathbf{y}$ and the squared Euclidean distance for $\mathcal{L}$ are commonly used (Lipman et al., 2024, Eq. 2.3). Albergo et al. (2023, Thm. 2.7) prove that the global minimum in equation 2.2 recovers $\bm{v}_{t}$ from equation 2.1. Other interpolants or loss functions (Song and Dhariwal, 2024; Kim et al., 2024) as well as penalizations facilitating one-step generation (Boffi et al., 2025, and references therein) have also been proposed.
@@
Pooladian et al. (2023) and Tong et al. (2024) (see also discussion in Lipman et al. 2024, §4.9.2) propose OT-FM, a much lighter approach that simply replaces $\pi_{I}$ by a coupling $\pi$ that should, ideally, approximate the OT coupling $\pi^{\star}=\arg\min_{\pi\in\Gamma(\mu_{0},\mu_{1})}\mathcal{C}(\pi)$. Compared to Reflow, which uses FM as an inner step, OT-FM adds a pre-processing effort to FM training, to select better noise/data pairs. Because the ground-truth OT coupling $\pi^{\star}$ is never known, however, they use OT matrices: sampling $n\approx 256$ points from both $\mu_{0}$ and $\mu_{1}$, they match them using an OT solver (e.g. the Hungarian algorithm, Kuhn 1955) and feed these pairs to the flow loss equation 2.2. Davtyan et al. (2025) proposed then LOOM-CFM, a hybrid approach that stores buffers of paired noise/data samples (using the Hungarian algorithm). Unfortunately, OT-FM yields only modest gains. Zhang et al. (2025) posit that this is due to the well-known statistical bias arising when approximating continuous OT couplings using samples (Hütter and Rigollet, 2021). To reduce the effect of this bias, they use significantly higher $n$ handled with a multi-GPU, multi-node implementation of the Sinkhorn algorithm, carefully ablating the role of the $\varepsilon$ regularization parameter. While encouraging, their results show better performance as $n$ grows and $\varepsilon$ shrinks to 0, an explosive cocktail of hyperparameters since the cost of running Sinkhorn (1964) grows as $O(n^{2}/\varepsilon^{2})$. Fundamentally, these approaches are bottlenecked by their reliance on the discrete view of OT problems – that is, they rely repeatedly on matching $n$ i.i.d. samples of noise and data.
@@
One can solve equation $\mathrm{P}_{\varepsilon}$ by solving the concave maximization semidual problem parameterized entirely by $\mathbf{g}$, that Genevay et al. (2016, Alg. 2) proposed to solve using averaged SGD (our Algorithm˜1)
@@
For any $\varepsilon\geq 0$, the values of equation $\mathrm{P}_{\varepsilon}$ and equation $\mathrm{S}_{\varepsilon}$ are equal. Moreover for $\varepsilon>0$, given a maximizing potential $\mathbf{g}^{\star}=(g^{\star}_{1},\dots,g^{\star}_{N})$ of equation $\mathrm{S}_{\varepsilon}$, one can recover a minimizer $\pi^{\star}_{\varepsilon}$ of equation $\mathrm{P}_{\varepsilon}$ that has density
@@ 3 Improved Optimization for Semidiscrete Optimal Transport
We start from the formulation presented by Genevay et al. (2016) and adapt it to the scales relevant to our work. We present new practical recommendations as well as a new theoretical analysis to inform the computation of SD-OT, with (or without) entropic regularization.
Marginal Estimation for Semidiscrete Couplings.
A crucial aspect of the primal-dual relationship in equation 2.5 is that it holds at optimality, but can also be used to create a coupling: one can plug a vector $\mathbf{g}\in{\mathbb{R}}^{N}$ into equation 2.5 to define a joint probability that only satisfies the first marginal constraint. The exponentiation in equation 2.5 can be rewritten using the softmax operator mapping a vector $\mathbf{z}=(z_{1},\dots,z_{N})$ to a probability vector in the simplex $\Delta^{N}$,
@@
By construction, the first marginal of $\pi_{\varepsilon,{\mathbf{g}}}$ is $\mu$: Indeed, the measure $\pi_{\varepsilon,{\mathbf{g}}}(\mathbf{x},\mathbf{y}_{j})$ in equation 3.2 sums to $\mathrm{d}\mu(\mathbf{x})$ when integrated against the data for a fixed $\mathbf{x}$, since summing w.r.t. index $j$ is akin to summing a weighted soft-max distribution. On the other end, $\pi_{\varepsilon,{\mathbf{g}}}$ would be an OT coupling if and only if its second marginal, defined as
$\mathbf{m}({\mathbf{g}}):=[\pi_{\varepsilon,{\mathbf{g}}}({\mathbb{R}}^{d},\mathbf{y}_{j})]_{j}\!=\!\int_{{\bm{x}}\in{\mathbb{R}}^{d}}\!\!\!\!\!\!\!\!\!\!\mathbf{s}_{\varepsilon,{\mathbf{g}}}(\mathbf{x})\mathrm{d}\mu(\mathbf{x})\in\Delta^{N}$ (3.3)
coincided with $\mathbf{b}$, by analogy to the Sinkhorn algorithm (Peyré and Cuturi, 2019, §4.14). Note that $\pi_{\varepsilon,{\mathbf{g}}}$ solves equation $\mathrm{P}_{\varepsilon}$ if $\mathbf{g}$} is a maximizer of equation $\mathrm{S}_{\varepsilon}$.
@@
When $\nu$ is uniform, by Lemma˜7, the above $\chi^{2}$ divergence is also equal to the rescaled squared norm of the semidual gradient. From equation 3.4, we propose in equation 3.5 an unbiased estimator using a batch of i.i.d. samples $\mathbf{x}_{1},\dots,\mathbf{x}_{B}\sim\mu$ that can be computed in $O(NB)$ time. Figure˜2 shows that we are able to effectively track ${\widehat{\chi}}^{2}$ as it decays towards zero along SGD updates.
@@
Suppose $c(\mathbf{x},\mathbf{y})=-\mathbf{x}^{T}\mathbf{y}$, $\delta\coloneqq\min_{j\neq j^{\prime}}\|\mathbf{y}_{j}-\mathbf{y}_{j^{\prime}}\|>0$, and $\mu$ admits a density w.r.t. the Lebesgue measure. Additionally, suppose the maximum $\mu$-surface area of any convex subset of ${\mathbb{R}}^{d}$ is bounded by $C^{\mathrm{max}}_{\mu}$, i.e. $\smallint_{\partial A}\mu(\mathbf{x})\mathrm{d}S(\mathbf{x})\leq C^{\mathrm{max}}_{\mu}$ for all convex $A\subseteq{\mathbb{R}}^{d}$, $S$ being the d1d-1-dimensional Euclidean surface measure. When $\mu=\mathcal{N}(0,{\bm{I}}_{d})$, $C^{\mathrm{max}}_{\mu}\leq 4d^{1/4}$ (Ball, 1993).
@@ 4 Semidiscrete transport for flow matching
Building on Section˜3, the OT problem is solved between Gaussian noise $\mu_{0}$ and the data measure $\mu_{1}$ supported on $(\mathbf{x}^{(1)}_{1},\dots,\mathbf{x}^{(N)}_{1})$. This framework can also accommodate the conditional generation setting, in which observations are paired with a $p$-dimensional condition, see Section˜2. We follow Zhang et al. (2025, §3) and use the dot-product cost $c(\mathbf{x},\mathbf{y}):=-\mathbf{x}^{T}\mathbf{y}$ (augmenting it with temperature $\beta$ for conditioning equation 2.3) and rescale $\varepsilon$ with the cost matrix standard deviation for a sampled reference batch. As illustrated in Figure 1, SD-FM works in two steps, (1) solve SD-OT between $\mu_{0}$ and $\mu_{1}$ before flow training, storing $\mathbf{g}^{\star}\in{\mathbb{R}}^{N}$ solving equation $\mathrm{S}_{\varepsilon}$; (2) in the FM train loop, pair a freshly sampled noise $\mathbf{x}_{0}$ to data $\mathbf{x}_{1}^{(k)}$, where $k\sim s_{\varepsilon,\bm{g}^{\star}}(\mathbf{x}_{0})\in\Delta^{N}$.
@@
Table 1: Complexity of I/OT/SD-FM given NN: dataset size, $d$: data dimension, $E$: FM training epochs, $n$: Sinkhorn batch size, $\varepsilon$: entropic regularization, $\Theta$: cost of computing FM loss’ gradient for a noise/data pair. $K$: SD SGD steps. OT-FM* is the cached variant of OT-FM. $\Theta$ dominates all costs, $N\geq n$, and $\varepsilon$ is small.
As recalled in equation 3.2, for $\varepsilon>0$, this amounts to a categorical sampling among $N$ points, while for $\varepsilon=0$ this reduces to an assignment to the $k^{\star}$-th point in the dataset, $k^{\star}=\arg\max_{k}g^{\star}_{k}+\langle\mathbf{x}_{0},\mathbf{x}^{(k)}_{1}\rangle$ (or sample uniformly among ties if they arise). The latter operation is slightly faster than categorical sampling, as observed in Fig. 4. The differences between these three approaches are highlighted in Algorithms 3, 4, 5, 6. In a nutshell, SD-FM uses a simple SGD precompute effort to fit $\mathbf{g}$, with a minor lookup effort to pair each noise with a datapoint, rather than making potentially costly calls to Sinkhorn in OT-FM. That lookup is a MIPS procedure (Shrivastava and Li, 2014) when $\varepsilon=0$, which gives room for further speed improvement when using data structures to bring down that $O(N)$ complexity.
@@
Note that the above is Tweedie’s formula since the flow matching velocity that transports $\rho_{0}$ to $\rho_{1}$ satisfies $(1-t){\bm{v}}_{t}(\mathbf{x})={\mathbb{E}}[\bm{X}_{1}-\bm{X}_{t}\,|\,\bm{X}_{t}=\mathbf{x}]$. While equation 4.1 relates score and velocity, the second equality above relies on the independence of $\bm{X}_{0}$ and $\bm{X}_{1}$; it may no longer hold under general couplings. Surprisingly, we find that for SD-OT couplings equation 4.1 still holds approximately for $\varepsilon\approx 0$.

### v2

@@ Abstract
Flow models parameterized as time-dependent velocity fields can generate data from noise by integrating an ODE. These models are often trained using flow matching, i.e. by sampling random pairs of noise and target points $(\mathbf{x}_{0},\mathbf{x}_{1})$ and ensuring that the velocity field is aligned, on average, with $\mathbf{x}_{1}-\mathbf{x}_{0}$ when evaluated along a time-indexed segment linking $\mathbf{x}_{0}$ to $\mathbf{x}_{1}$. While these noise/data pairs are sampled independently by default, they can also be selected more carefully by matching batches of $n$ noise to $n$ target points using an optimal transport (OT) solver. Although promising in theory, the OT flow matching (OT-FM) approach (Pooladian et al., 2023; Tong et al., 2024) is not widely used in practice. Zhang et al. (2025) pointed out recently that OT-FM truly starts paying off when the batch size $n$ grows significantly, which only a multi-GPU implementation of the Sinkhorn algorithm can handle. Unfortunately, the pre-compute costs of running Sinkhorn can balloon, requiring $O(n^{2}/\varepsilon^{2})$ operations for every $n$ pairs used to fit the velocity field, where $\varepsilon$ is a regularization parameter that should be typically small to yield better results. To fulfill the theoretical promises of OT-FM, we propose to move away from batch-OT and rely instead on a semidiscrete formulation that can leverage the fact that the target dataset is usually of finite size $N$. The SD-OT problem is solved by estimating a dual potential vector of size $N$ using SGD; using that vector, freshly sampled noise vectors at train time can then be matched with data points at the cost of a maximum inner product search (MIPS) over the dataset with a cost $O(N)$. Semidiscrete FM (SD-FM) removes the quadratic dependency on $n/\varepsilon$ that bottlenecks OT-FM. SD-FM beats both FM and OT-FM on all training metrics and inference budget constraints, across multiple datasets, on unconditional/conditional generation, or when using mean-flow models.
@@ 1 Introduction
Figure 1: I-FM (left) assigns noise to data purely at random. OT-FM (middle-left) samples batches of $n$ noise and $n$ data points and re-aligns them with an optimal matching permutation $\sigma^{\star}$. These matches are, however, inherently unstable, as $n$ points do not reflect the whole noise distribution nor the dataset. Increasing drastically $n$ can mitigate this issue (Zhang et al., 2025), but at a significant cost. Our method, SD-FM (right), solves these issues in two steps: in a precompute phase, the semidiscrete OT problem (parameterized as a vector of size NN, the dataset size) is solved using SGD. At FM train time , each newly sampled noise is assigned to a data point using a maximum inner product search, Laguerre cells (Mérigot, 2011) being illustrated in the plot. Our figure uses no entropic regularization ($\varepsilon=0$) and a neg-dot-product cost for simplicity, see (2.4) for more generality.
@@ 2 Background and related work
Access to $\bm{v}_{t}$ then allows for samples $\bm{X}_{0}\sim\mu_{0}$ to be transformed into samples $\bm{X}_{1}\sim\mu_{1}$ through evolution of (2.1). Training a neural approximation $\bm{v}_{\theta}$ for the flow field $\bm{v}$ can be done by sampling $(\bm{X}_{0},\bm{X}_{1})\sim\pi$ and $t\in[0,1]$, setting $\bm{X}_{t}:=\varphi_{t}(\bm{X}_{0},\bm{X}_{1})$ and minimizing
@@
The linear interpolant $\varphi_{t}(\mathbf{x},\mathbf{y}):=(1-t)\mathbf{x}+t\mathbf{y}$ and the squared Euclidean distance for $\mathcal{L}$ are commonly used (Lipman et al., 2024, Eq. 2.3). Albergo et al. (2023, Thm. 2.7) prove that the global minimum in (2.2) recovers $\bm{v}_{t}$ from (2.1). Other interpolants or loss functions (Song and Dhariwal, 2024; Kim et al., 2024) as well as penalizations facilitating one-step generation (Boffi et al., 2025, and references therein) have also been proposed.
@@
Pooladian et al. (2023) and Tong et al. (2024) (see also discussion in Lipman et al. 2024, §4.9.2) propose OT-FM, a much lighter approach that simply replaces $\pi_{I}$ by a coupling $\pi$ that should, ideally, approximate the OT coupling $\pi^{\star}=\arg\min_{\pi\in\Gamma(\mu_{0},\mu_{1})}\mathcal{C}(\pi)$. Compared to Reflow, which uses FM as an inner step, OT-FM adds a pre-processing effort to FM training, to select better noise/data pairs. Because the ground-truth OT coupling $\pi^{\star}$ is never known, however, they use OT matrices: sampling $n\approx 256$ points from both $\mu_{0}$ and $\mu_{1}$, they match them using an OT solver (e.g. the Hungarian algorithm, Kuhn 1955) and feed these pairs to the flow loss (2.2). Davtyan et al. (2025) proposed then LOOM-CFM, a hybrid approach that stores buffers of paired noise/data samples (using the Hungarian algorithm). Unfortunately, OT-FM yields only modest gains. Zhang et al. (2025) posit that this is due to the well-known statistical bias arising when approximating continuous OT couplings using samples (Hütter and Rigollet, 2021). To reduce the effect of this bias, they use significantly higher $n$ handled with a multi-GPU, multi-node implementation of the Sinkhorn algorithm, carefully ablating the role of the $\varepsilon$ regularization parameter. While encouraging, their results show better performance as $n$ grows and $\varepsilon$ shrinks to 0, an explosive cocktail of hyperparameters since the cost of running Sinkhorn (1964) grows as $O(n^{2}/\varepsilon^{2})$. Fundamentally, these approaches are bottlenecked by their reliance on the discrete view of OT problems – that is, they rely repeatedly on matching $n$ i.i.d. samples of noise and data.
@@
One can solve ($\mathrm{P}_{\varepsilon}$) by solving the concave maximization semidual problem parameterized entirely by $\mathbf{g}$, that Genevay et al. (2016, Alg. 2) proposed to solve using averaged SGD (our Algorithm˜1)
@@
For any $\varepsilon\geq 0$, the values of ($\mathrm{P}_{\varepsilon}$) and ($\mathrm{S}_{\varepsilon}$) are equal. Moreover for $\varepsilon>0$, given a maximizing potential $\mathbf{g}^{\star}=(g^{\star}_{1},\dots,g^{\star}_{N})$ of ($\mathrm{S}_{\varepsilon}$), one can recover a minimizer $\pi^{\star}_{\varepsilon}$ of ($\mathrm{P}_{\varepsilon}$) that has density
@@ 3 Improved Optimization for Semidiscrete Optimal Transport
We adapt the approach of Genevay et al. (2016) to the much larger scales used in this work through two novel contributions: we propose an unbiased convergence criterion to monitor convergence, and a theoretical analysis to inform the computation of SD-OT that works also for the case $\varepsilon=0$. Genevay et al. (2016) had not considered that case, which turns out to be the most promising when using SD-OT within the SD-FM methodology presented next in §4.
Marginal Estimation for SD Couplings.
The primal-dual relationship in (2.5) holds at optimality, but can also be used to create a coupling by plugging a vector $\mathbf{g}\in{\mathbb{R}}^{N}$ into (2.5) to define a joint probability that only satisfies the first marginal constraint. The exponentiation in (2.5) can be rewritten using the softmax operator mapping a vector $\mathbf{z}=(z_{1},\dots,z_{N})$ to a vector in the simplex $\Delta^{N}$,
@@
By construction, the first marginal of $\pi_{\varepsilon,{\mathbf{g}}}$ is $\mu$: Indeed, the measure $\pi_{\varepsilon,{\mathbf{g}}}(\mathbf{x},\mathbf{y}_{j})$ in (3.2) sums to $\mathrm{d}\mu(\mathbf{x})$ when integrated against the data for a fixed $\mathbf{x}$, since summing w.r.t. index $j$ is akin to summing a weighted soft-max distribution. On the other end, $\pi_{\varepsilon,{\mathbf{g}}}$ would be an OT coupling if and only if its second marginal, defined as
$\mathbf{m}({\mathbf{g}}):=[\pi_{\varepsilon,{\mathbf{g}}}({\mathbb{R}}^{d},\mathbf{y}_{j})]_{j}\!=\!\int_{{\bm{x}}\in{\mathbb{R}}^{d}}\!\!\!\!\!\!\!\!\!\!\mathbf{s}_{\varepsilon,{\mathbf{g}}}(\mathbf{x})\mathrm{d}\mu(\mathbf{x})\in\Delta^{N}.$ (3.3)
coincided with $\mathbf{b}$, by analogy to the Sinkhorn algorithm (Peyré and Cuturi, 2019, §4.14). Note that $\pi_{\varepsilon,{\mathbf{g}}}$ solves ($\mathrm{P}_{\varepsilon}$) if $\mathbf{g}$} is a maximizer of ($\mathrm{S}_{\varepsilon}$).
@@
When $\nu$ is uniform, by Lemma˜7, the above $\chi^{2}$ divergence is also equal to the rescaled squared norm of the semidual gradient. From (3.4), we propose in (3.5) an unbiased estimator using a batch of i.i.d. samples $\mathbf{x}_{1},\dots,\mathbf{x}_{B}\sim\mu$ that can be computed in $O(NB)$ time. Figure˜2 shows that we are able to effectively track ${\widehat{\chi}}^{2}$ as it decays towards zero along SGD updates.
@@
Suppose $c(\mathbf{x},\mathbf{y})=-\mathbf{x}^{\top}\mathbf{y}$, $\delta\coloneqq\min_{j\neq j^{\prime}}\|\mathbf{y}_{j}-\mathbf{y}_{j^{\prime}}\|>0$, and $\mu$ admits a density w.r.t. the Lebesgue measure. Additionally, suppose the maximum $\mu$-surface area of any convex subset of ${\mathbb{R}}^{d}$ is bounded by $C^{\mathrm{max}}_{\mu}$, i.e. $\smallint_{\partial A}\mu(\mathbf{x})\mathrm{d}S(\mathbf{x})\leq C^{\mathrm{max}}_{\mu}$ for all convex $A\subseteq{\mathbb{R}}^{d}$, $S$ being the d1d-1-dimensional Euclidean surface measure. When $\mu=\mathcal{N}(0,{\bm{I}}_{d})$, $C^{\mathrm{max}}_{\mu}\leq 4d^{1/4}$ (Ball, 1993).
@@ 4 Semidiscrete transport for flow matching
Building on Section˜3, the OT problem is solved between Gaussian noise $\mu_{0}$ and the data measure $\mu_{1}$ supported on $(\mathbf{x}^{(1)}_{1},\dots,\mathbf{x}^{(N)}_{1})$. This framework can also accommodate the conditional generation setting, in which observations are paired with a $p$-dimensional condition, see Section˜2. We follow Zhang et al. (2025, §3) and use the dot-product cost $c(\mathbf{x},\mathbf{y}):=-\mathbf{x}^{T}\mathbf{y}$ (augmenting it with temperature $\beta$ for conditioning (2.3)) and rescale $\varepsilon$ with the cost matrix standard deviation for a sampled reference batch. As illustrated in Figure 1, SD-FM works in two steps, (1) solve SD-OT between $\mu_{0}$ and $\mu_{1}$ before flow training, storing $\mathbf{g}^{\star}\in{\mathbb{R}}^{N}$ solving ($\mathrm{S}_{\varepsilon}$); (2) in the FM train loop, pair a freshly sampled noise $\mathbf{x}_{0}$ to data $\mathbf{x}_{1}^{(k)}$, where $k\sim s_{\varepsilon,\bm{g}^{\star}}(\mathbf{x}_{0})\in\Delta^{N}$.
@@
Table 1: Complexity of I/OT/SD-FM given NN: dataset size, $d$: data dimension, $E$: FM training epochs, $n$: Sinkhorn batch size, $\varepsilon$: entropic regularization, $\Theta$: cost of computing FM loss’ gradient for a pair. $K$: SD SGD steps. OT-FM* is the cached variant of OT-FM. $\Theta$ dominates all costs, $N\geq n$, $\varepsilon\approx 0$.
As recalled in (3.2), for $\varepsilon>0$, this amounts to a categorical sampling among $N$ points, while for $\varepsilon=0$ this reduces to an assignment to the $k^{\star}$-th point in the dataset, $k^{\star}=\arg\max_{k}g^{\star}_{k}+\langle\mathbf{x}_{0},\mathbf{x}^{(k)}_{1}\rangle$ (or sample uniformly among ties if they arise). The latter operation is slightly faster than categorical sampling, as observed in Fig. 4. The differences between these three approaches are highlighted in Algorithms 3, 4, 5, 6. In a nutshell, SD-FM uses a simple SGD precompute effort to fit $\mathbf{g}$, with a minor lookup effort to pair each noise with a datapoint, rather than making potentially costly calls to Sinkhorn in OT-FM. That lookup is a MIPS procedure (Shrivastava and Li, 2014) when $\varepsilon=0$, solved exactly, incurring a $Nd$ cost (noise $\times$ data dot-products) and a max over $N$ values. This might be sped-up with the use of approximate MIPS, left for future work.
@@
Note that the above is Tweedie’s formula since the flow matching velocity that transports $\rho_{0}$ to $\rho_{1}$ satisfies $(1-t){\bm{v}}_{t}(\mathbf{x})={\mathbb{E}}[\bm{X}_{1}-\bm{X}_{t}\,|\,\bm{X}_{t}=\mathbf{x}]$. While (4.1) relates score and velocity, the second equality above relies on the independence of $\bm{X}_{0}$ and $\bm{X}_{1}$; it may no longer hold under general couplings. Surprisingly, we find that for SD-OT couplings (4.1) still holds approximately for $\varepsilon\approx 0$.